SoCuteBleh:

The Baldwin family walked four over five of a mile in four over nine of an hour. What is their unit rate in miles per hour? (5 points)

Mercury:

unit rate = # of miles / # of hours = \[\frac{ \frac{ 4 }{ 5 } }{ \frac{ 4 }{ 9 } }\] to evaluate this, flip the second fraction and change the division sign between the fractions, to a multiplication sign. then from there, it's simple fraction multiplication, where you multiply the numerators, multiply the denominators, and re-write the final fraction